# Wave-particle correlations in multiphoton resonances of coherent light-matter interaction

### 07 Mar 2024 v2

@@ Abstract
Wave/particle duality, photon blockade, Jaynes-Cummings model, direct photodetection, balanced homodyne detection, squeezed light, quantum interference
@@ I Introduction
Our narrative is structured as follows. In Sec. II, we formulate the driven dissipative unconditional dynamical description of an ensemble-averaged evolution. Section III is devoted to the derivation of an analytical expression for the waiting-time distribution of the forwards-scattered photons in the limit of vanishing spontaneous emission for the two-state atom – the so-called “zero system-size” limit of absorptive optical bistability – which is of special interest to the persistence of photon blockade Carmichael (2015); Mavrogordatos (2019). In Sec. IV, we use the minimal four-state model of a cascaded two-photon resonance to derive analytical results for the transmission spectrum and for the spectrum of squeezing. Finally, in Sec. V, we detail the unraveling scheme based on the wave-particle correlator applied to the light emanating from the JC oscillator under multiphoton resonance operation, before summarizing our main findings in the Conclusions.
@@ II The source master equation: dressing of dressed states in action
Figure 2: Conditioned separation vs. unconditional coexistence of timescales in the particle aspect of light. (a) Sample trajectory of the conditioned intracavity photon number $\braket{a^{\dagger}a(t)}_{\rm REC}$ from the decomposition of the ME (1) under direct photodetection ($r=1$) with initial condition $\rho(0)=|1,-\rangle\langle 1,-|$. The green (red crossed) strokes underneath indicate spontaneous (cavity) emission events. The surface plots in frame (a) depict the transient conditioned Wigner function of the cavity field from the high to the lower photon excitation across the two cavity emissions marked A and B, while the contour plots in both frames depict the unconditioned Wigner function from the steady-state numerical solution of the ME (1) in MATLAB’s Quantum Optics Toolbox Tan (1999). (b) Steady-state waiting-time distribution of the forwards-scattered light, $W_{\rm ss}(\tau)\equiv\kappa^{-1}w_{{\rm ss},\rightarrow}(\tau)$ from Eq. (6), analytically derived from the minimal model and plotted for two different values of the driving strength such that: $\Omega/\kappa=10/\sqrt{2}$ for (i) [corresponding to the same drive strength as in (a)] and $1/\sqrt{2}$ for (ii). The lower inset depicts the numerically determined $W_{\rm ss}(\tau)$ from the quantum regression formula, alongside the superposition of the quantum beat to the first half period of the Rabi oscillation. In frame [(a), (b)], $\gamma/(2\kappa)=[1,0]$, while in both frames $g/\kappa=1,000$. For (a), a Monte-Carlo (MC) algorithm was developed ad hoc in MATLAB using a basis truncated at the 14-photon level.
@@
The intermediate state are not to be discarded since they take part in the cascaded process. The shifted energy levels dressed by the drive as second-order corrections in $\varepsilon_{d}/g$ are
@@
while the effective two-photon Rabi frequency is $\Omega=2\sqrt{2}\,\varepsilon_{d}^{2}/g$ Mavrogordatos and Lledó (2021). In the effective ME (2), we define $\mathcal{D}[X](\rho)\equiv X\rho X^{\dagger}-(1/2)\{X^{\dagger}X,\rho\}$, while the transition rates between the four levels – in the special case where $\gamma=2\kappa$ – are Shamailov et al. (2010)
@@
Three timescales revealed by the minimal model are present in the unraveling of Fig. 2(a) under direct photodetection as well as in the numerically obtained waiting-time distribution of Fig.Fig. 2(a). The fastest of these three corresponds to the quantum beat, with frequency equal to $\nu\equiv(\tilde{E}_{2}-\tilde{E}_{1})/\hbar=2g+\delta_{2}-\delta_{1}=2g+\mathcal{O}(\varepsilon_{d}^{2}/g)$. An intermediate timescale, which is missed in the secular approximation and therefore in the analytical treatment, is present in the numerically-extracted evolution due to the asymmetry between the excitation paths $|\xi_{3}\rangle\to|\xi_{1}\rangle\to|\xi_{0}\rangle$ and $|\xi_{3}\rangle\to|\xi_{2}\rangle\to|\xi_{0}\rangle$, since $\Gamma_{31}/\Gamma_{32}\approx 5.8$ for $\gamma/(2\kappa)=0$ (see Sec. 3.2 of Shamailov et al. (2010) and, e.g., the field-amplitude oscillations in the top panel of Fig. 6). In a frame rotating with the drive and for $\Delta\omega_{d}=-g/\sqrt{2}$, the unperturbed energies of the intermediate states are $E^{(0)}_{1,2}=g(1\mp\sqrt{2})/\sqrt{2}$ whereas the outer states $|\xi_{3}\rangle$ and $|\xi_{0}\rangle$ have zero energy. It follows that the excitation path $|\xi_{3}\rangle\to|\xi_{1}\rangle\to|\xi_{0}\rangle$ is associated with a periodic structure including $2\sqrt{2}/(\sqrt{2}-1)\approx 6.8$ quantum beat cycles while the path $|\xi_{3}\rangle\to|\xi_{2}\rangle\to|\xi_{0}\rangle$ is associated with a period of about $1.2$ times the beat cycle, manifested as an occasional deviation from the pattern dictated by the dominant path. The slowest timescale corresponds to the semiclassical oscillation with frequency $2\Omega=4\sqrt{2}\varepsilon_{d}^{2}/g$ Mavrogordatos and Lledó (2021) and is associated with driving a saturable (effective) two-level transition, similar to ordinary resonance fluorescence.
@@ III Particle aspect: intensity correlation and waiting-time distribution
Let us keep following further the trajectory generated by a direct-photodetection unraveling of the ME. The emission of the first photon from the second pair reinitiates the semiclassical oscillation, but now one where the superimposed quantum beat does not feature as prominently. This is due to the weaker participation of the states $|\xi_{1}\rangle$ and $|\xi_{2}\rangle$ in the conditioned state resulting after the second spontaneous emission event. The second recorded photon emission in the forwards direction follows an even-state superposition in the cavity-field distribution, which is again reduced to a conditioned state with the Wigner function of a SPACS. Both quantum-superposition states in the distributions at the two ends of the jump are framed by remnants of steady-state bimodality.
Figure 4: Photon blockade and intensity correlation functions. (a) Steady-state intracavity photon number obtained from the numerical solution to the ME (1), plotted against the (absolute value of the) scaled drive detuning $|\Delta\omega_{d}|/g$ for three increasing values of the drive strength $\varepsilon_{d}/g$: $0.03,0.055$ and $0.16$ in one-to-one correspondence to the three resonance sequences of growing excitation in the cavity. Intersection 1 corresponds to the operation point selected for Fig. 5(a), 2 to Figs. 5(b, c), and 3 to Fig. 5(d). The three interaction points correspond to the inputs for the wave-particle correlator whose outputs are depicted in Fig. 5. The steady-state intensity correlation function of the forwards-scattered photons $g_{\rm ss}^{(2)}(\tau)$ corresponding to these three operation points are plotted in frames (b), (c), (d) respectively. The two insets in frames (b), (c) depict the unlabeled spectrum of squeezing $\mathcal{S}_{\rightarrow}^{\theta}(\omega)$ against $(\omega-\omega_{0})/\kappa$, calculated from Eq. (44) and the same conditions as in Figs. 5(a) [for $\theta=3\pi/4$] and (c) [for $\theta=\pi/4$], respectively . The inset in frame (d) focuses on the initial repetitive pattern formed by the superposition of quantum beats corresponding to Fig. 5(d). The intensity correlation functions have been numerically calculated in MATLAB’s Quantum Optics Toolbox Tan (1999) using the quantum regression formula Carmichael (1999) in a Hilbert space of 14 photon levels.
To assess the statistics of the light field stemming from its corpuscular nature, we have derived the waiting-time distribution of the forwards-scattered radiation Carmichael et al. (1989); Car (1993); Brandes (2008) as (see Appendix A),
@@
an expression plotted in Fig. 2(b). The time $\tau$ waited between successive photoelectric emissions detected with perfect efficiency follows the superposition of the quantum beat [term $\propto\cos(\nu\tau)$] onto Rabi oscillations, the latter being developed as a characteristic ringing for increasing drive strength [last term in Eq. (6)]. In contrast to ordinary resonance fluorescence, the distribution minima do not reach zero but are rather bound by the asymptote $\kappa\frac{1}{3}e^{-\kappa\tau}$ defined by the occupation of the intermediate levels in the cascaded process. The higher bound constraining the superposition of oscillations is reached in the limit $\varepsilon_{d}/\kappa\gg 1$. Figure 2(b) shows a very good agreement between Eq. (6) and the numerical solution of the ME (1) combined with applying the quantum regression formula.
@@ IV Wave aspect: the spectrum of squeezing
Setting $r=0$ cancels the triggering in the wave-particle correlator and actuates solely the BHD arm, typically used for the measurement of amplitude-squeezing based on the ability of light to interfere with a strong coherent LO field. These signatures speak of the continuous (wave) character of light. The spectrum of squeezing can be calculated by constructing the photocurrent autocorrelation $\overline{i(t)i(t+\tau)}$ for the selected quadrature of the field $A_{\theta}\equiv\frac{1}{2}(ae^{-i\theta}+a^{\dagger}e^{i\theta})$ as a time average over ergodic records Cresser (2001), and taking the Fourier transform Carmichael (2008).
@@
Let us now revert to unconditional dynamics obeying the master equation (1) with reference to the nonclassical attributes of the particle and wave aspects separately, each revealed by a suitable multi-time correlation function. Figure 4 presents the complementary nonexclusive probability densities expressed through the normalized second-order correlation function Car (1993) for the forwards scattered field. We witness the transition from photon bunching to antibunching as steady-state bimodality builds up alongside a more pronounced Rabi splitting. Consequences of this change in photon statistics are revealed in the individual quantum trajectories of Figs. 5(a ,b), where the two occurring photon emissions are being spaced further apart in the course of an unraveling which combines both particle and wave aspect of light. When we move away from the two-photon resonance peak we come across a superposition of quantum beats with different frequencies and smaller amplitudes [see Fig 4(d)], a consequence of significant detuning of the drive field from an effective two-level resonance structure formed between the dressed JC eigenstates.
@@ V Particles triggering the sampling of waves
At this stage, let us put both arms in operation by setting $r=1/2$ to produce a sampling of a nonlinear stochastic Schrödinger evolution solved by the conditioned wavefunction $|\psi_{c}(t)\rangle$. Our interest is with the full function of the wave-particle correlator, with reference to Fig. 1, in its ability to jointly detect the dual aspects of the incoming radiation. The set $(r,\theta)$ determines each one of the infinite unravelings of the ME: $10<r<$ 1, while the phase $\theta$ of the local oscillator selects the quadrature of interest. The negative spectrum of squeezing we previously depicted for the two-photon resonance driving of the JC oscillator signifies a redistribution of fluctuations between the different quadratures of the cavity field. Here we aim to analyze the quantum fluctuations of the field as ostensible deviations from the steady state, combining conditional measurements with the process of quantum measurement.
We write the master equation for the source – a driven JC oscillator in its zero “system size limit” $\gamma/(2\kappa)\to 0$ – extended now to include the local oscillator mode tuned to the frequency of the coherent drive $\omega_{d}$. The density matrix $\tilde{\rho}$ of the composite system in a frame rotating with $\omega_{d}$ reads Reiner et al. (2001):
@@
where $c$ and $c^{\dagger}$ are the raising and lowering operators for the local oscillator driven mode with excitation amplitude $\beta$. Under the BHD scheme of Fig. 1, the total measured fields at the photodetectors 1 and 2 (in photon flux units) read
@@
Under the assumption that the local oscillator mode is unaffected by the signal mode, the density matrix factorizes into a piece corresponding to the driven JC oscillator under photon blockade, and a piece corresponding to the local oscillator, $\tilde{\rho}=\tilde{\rho}_{s}|\beta\rangle\langle\beta|$. Defining the local oscillator flux as $f\equiv\kappa_{\rm LO}|\beta|^{2}$, we construct the following collapse super-operators,
@@
There are two types of detections which take place under the conditioned balanced homodyne detection (BHD) measurement scheme. The first comes from Eq. (10b) and corresponds to detection “clicks” at the avalanche photodiode (APD). These collapses occur with probability $2\kappa r\langle\psi_{c}(t)|a^{\dagger}a|\psi_{c}(t)\rangle dt$. The second type of collapses concerns the “clicks” registered through the BHD. In any realistic homodyne measurement the local oscillator photon flux $f$ is many orders of magnitude larger than the signal flux, here $2\kappa(1-r)\langle\psi_{c}(t)|a^{\dagger}a|\psi_{c}(t)\rangle$. This means that under the action of (8), a photoelectric emission corresponds with high probability to an annihilation of a local oscillator photon. There is only a small probability ${}f/[2\kappa\langle\psi_{c}(t)|a^{\dagger}a|\psi_{c}(t)\rangle]$ that a photon originated from the JC oscillator. We should note here that these two possibilities exist as a superposition and not as a classical choice, either one or the other. Despite the fact that the collapses are very small, on the characteristic timescale $(2\kappa)^{-1}$ for fluctuations in the signal field they occur very often. The quantum mapping into a stochastic differential equation – a Schrödinger equation with a stochastic non-Hermitian Hamiltonian – involves a coarse graining in time, an expansion of the non-unitary evolution in powers of $\sqrt{2\kappa(1-r)/f}$ and the consideration of a stochastic process for the number of the emitted photoelectrons which depends on a conditioned wavefunction that satisfies the very same Schrödinger equation. Taking that path leads to the following expressions for the difference in photocurrents between detectors 1 and 2 (in Fig. 1), and the wavefunction propagation between photodetections at the APD:
@@
In Eqs. (12) and (13), $\langle A_{\theta}(t)\rangle_{c}$ is the conditioned average $\langle A_{\theta}(t)\rangle_{c}\equiv\tfrac{1}{2}\langle\psi_{c}(t)|e^{i\theta}a^{\dagger}+ae^{-i\theta}|\psi(t)\rangle$, calculated with the normalized conditioned wavefunction $|\psi_{c}(t)\rangle$, $B$ is the detection bandwidth and $dW_{t}$ is the Wiener noise increment, the same in both equations (12), (13). In the Monte-Carlo algorithm developed to simulate the wave-particle correlator unravelling, the stochastic differential equation (13) is solved by means of an explicit order 2.0 weak scheme proposed by Kloeden and Platen Kloeden and Platen (1999).
Figure 5: Sampling a homodyne current under low-amplitude nonlinearity. Individual trajectories of the conditioned intracavity photon number (positive-valued plots) and the conditioned field (plots with alternating sign), measured for $r=0.5$ with initial condition $\rho(0)=|1,-\rangle\langle 1,-|$ at different values of the driving frequency and strength $(\Delta\omega_{d}/g,\varepsilon_{d}/g)$: (a) $(-0.7114,0.03)$, (b, c) $(-0.7114,0.055)$ and (d) $(0.545,0.16)$. The inset contour plots in each frame depict the steady-state Wigner function, while the white dashed line indicates the direction set by the phase $\theta$ of the LO. The two additional insets in frame (d) depict the variation of the correlation function of the forwards scattered photons $g_{\rm ss}^{(2)}(\tau)$ over an average photon lifetime, at the peaks of the three ($n=3$) and four ($n=4$) photon resonances for $\varepsilon_{d}/g=0.16$. The shaded insets of (b, c) depict unconditional wave-particle correlation functions $h_{\pi/4}(\tau)$ and $h_{3\pi/4}(\tau)$, respectively, normalized to unity at long delays. In all frames, $g/\kappa=200$ and $\gamma/(2\kappa)=0$. A MC algorithm unravelling the ME (1) under the action of the wave-particle correlator was developed ad hoc in MATLAB using a basis truncated at the 14-photon level Kloeden and Platen (1999).

### 13 Mar 2024 v3

@@ Abstract
Wave/particle duality, photon blockade, Jaynes-Cummings model, direct photodetection, waiting-time distribution, balanced homodyne detection, squeezed light, quantum interference
@@ I Introduction
Our narrative is structured as follows. In Sec. II, we formulate the driven dissipative unconditional dynamical description of an ensemble-averaged evolution. Section III is devoted to the derivation of an analytical expression for the waiting-time distribution of the forward-scattered photons in the limit of vanishing spontaneous emission for the two-state atom – the so-called “zero system-size” limit of absorptive optical bistability – which is of special interest to the persistence of photon blockade Carmichael (2015); Mavrogordatos (2019). In Sec. IV, we use the minimal four-state model of a cascaded two-photon resonance to derive analytical results for the transmission spectrum and for the spectrum of squeezing. Finally, in Sec. V, we detail the unraveling scheme based on the wave-particle correlator applied to the light emanating from the JC oscillator under multiphoton resonance operation, before summarizing our main findings in the Conclusions.
@@ II The source master equation: dressing of dressed states in action
Figure 2: Conditioned separation vs. unconditional coexistence of timescales in the particle aspect of light. (a) Sample trajectory of the conditioned intracavity photon number $\braket{a^{\dagger}a(t)}_{\rm REC}$ from the decomposition of the ME (1) under direct photodetection ($r=1$) with initial condition $\rho(0)=|1,-\rangle\langle 1,-|$. The green (red crossed) strokes underneath indicate spontaneous (cavity) emission events. The surface plots in frame (a) depict the transient conditioned Wigner function of the cavity field from the high to the lower photon excitation across the two cavity emissions marked A and B, while the contour plots in both frames depict the unconditioned Wigner function from the steady-state numerical solution of the ME (1) in Matlab’s Quantum Optics Toolbox Tan (1999). (b) Steady-state waiting-time distribution of the forward-scattered light, $W_{\rm ss}(\tau)\equiv\kappa^{-1}w_{{\rm ss},\rightarrow}(\tau)$ from Eq. (6), analytically derived from the minimal model and plotted for two different values of the driving strength such that: $\Omega/\kappa=10/\sqrt{2}$ for (i) [corresponding to the same drive strength as in (a)] and $1/\sqrt{2}$ for (ii). The lower inset depicts the numerically determined $W_{\rm ss}(\tau)$ from the quantum regression formula, alongside the superposition of the quantum beat to the first half period of the Rabi oscillation. In frame [(a), (b)], $\gamma/(2\kappa)=[1,0]$, while in both frames $g/\kappa=1,000$. For (a), a Monte-Carlo (MC) algorithm was developed ad hoc in Matlab using a basis truncated at the 14-photon level.
@@
The intermediate states are not to be discarded since they take part in the cascaded process. The shifted energy levels dressed by the drive as second-order corrections in $\varepsilon_{d}/g$ are
@@
while the effective two-photon Rabi frequency is $\Omega=2\sqrt{2}\,\varepsilon_{d}^{2}/g$ Mavrogordatos and Lledó (2021). In the effective ME (2), we define $\mathcal{D}[X](\rho)\equiv X\rho X^{\dagger}-(1/2)\{X^{\dagger}X,\rho\}$, while the transition rates between the four levels, in the special case where $\gamma=2\kappa$, are Shamailov et al. (2010)
@@
Three timescales revealed by the minimal model are present in the unraveling of Fig. 2(a) under direct photodetection as well as in the numerically obtained waiting-time distribution of Fig. 2(a). The fastest of these three corresponds to the quantum beat, with frequency equal to $\nu\equiv(\tilde{E}_{2}-\tilde{E}_{1})/\hbar=2g+\delta_{2}-\delta_{1}=2g+\mathcal{O}(\varepsilon_{d}^{2}/g)$. An intermediate timescale, which is missed in the secular approximation and therefore in the analytical treatment, is present in the numerically-extracted evolution due to the asymmetry between the excitation paths $|\xi_{3}\rangle\to|\xi_{1}\rangle\to|\xi_{0}\rangle$ and $|\xi_{3}\rangle\to|\xi_{2}\rangle\to|\xi_{0}\rangle$, since $\Gamma_{31}/\Gamma_{32}\approx 5.8$ for $\gamma/(2\kappa)=0$ (see Sec. 3.2 of Shamailov et al. (2010) and, e.g., the field-amplitude oscillations in the top panel of Fig. 6). In a frame rotating with the drive and for $\Delta\omega_{d}=-g/\sqrt{2}$, the unperturbed energies of the intermediate states are $E^{(0)}_{1,2}=g(1\mp\sqrt{2})/\sqrt{2}$ whereas the outer states $|\xi_{3}\rangle$ and $|\xi_{0}\rangle$ have zero energy. It follows that the excitation path $|\xi_{3}\rangle\to|\xi_{1}\rangle\to|\xi_{0}\rangle$ is associated with a periodic structure including $2\sqrt{2}/(\sqrt{2}-1)\approx 6.8$ quantum beat cycles while the path $|\xi_{3}\rangle\to|\xi_{2}\rangle\to|\xi_{0}\rangle$ is associated with a period of about $1.2$ times the beat cycle, manifested as an occasional deviation from the pattern dictated by the dominant path. The slowest timescale corresponds to the semiclassical oscillation with frequency $2\Omega=4\sqrt{2}\varepsilon_{d}^{2}/g$ Mavrogordatos and Lledó (2021) and is associated with driving a saturable (effective) two-level transition, similar to ordinary resonance fluorescence.
@@ III Particle aspect: intensity correlation and waiting-time distribution
Let us keep following further the trajectory generated by a direct-photodetection unraveling of the ME. The emission of the first photon from the second pair reinitiates the semiclassical oscillation, but now one where the superimposed quantum beat does not feature as prominently. This is due to the weaker participation of the states $|\xi_{1}\rangle$ and $|\xi_{2}\rangle$ in the conditioned state resulting after the second spontaneous emission event. The second recorded photon emission in the forward direction follows an even-state superposition in the cavity-field distribution, which is again reduced to a conditioned state with the Wigner function of a SPACS. Both quantum-superposition states in the distributions at the two ends of the jump are framed by remnants of steady-state bimodality.
Figure 4: Photon blockade and intensity correlation functions. (a) Steady-state intracavity photon number obtained from the numerical solution to the ME (1), plotted against the (absolute value of the) scaled drive detuning $|\Delta\omega_{d}|/g$ for three increasing values of the drive strength $\varepsilon_{d}/g$: $0.03,0.055$ and $0.16$ in one-to-one correspondence to the three resonance sequences of growing excitation in the cavity. Intersection 1 corresponds to the operation point selected for Fig. 5(a), 2 to Figs. 5(b, c), and 3 to Fig. 5(d). The three interaction points correspond to the inputs for the wave-particle correlator whose outputs are depicted in Fig. 5. The steady-state intensity correlation function of the forward-scattered photons $g_{\rm ss}^{(2)}(\tau)$ corresponding to these three operation points are plotted in frames (b), (c), (d) respectively. The two insets in frames (b), (c) depict the unlabeled spectrum of squeezing $\mathcal{S}_{\rightarrow}^{\theta}(\omega)$ against $(\omega-\omega_{0})/\kappa$, calculated from Eq. (44) and the same conditions as in Figs. 5(a) [for $\theta=3\pi/4$] and (c) [for $\theta=\pi/4$], respectively . The inset in frame (d) focuses on the initial repetitive pattern formed by the superposition of quantum beats corresponding to Fig. 5(d). The intensity correlation functions have been numerically calculated in Matlab’s Quantum Optics Toolbox Tan (1999) using the quantum regression formula Carmichael (1999) in a Hilbert space of 14 photon levels.
To assess the statistics of the light field stemming from its corpuscular nature, we have derived the waiting-time distribution of the forward-scattered radiation Carmichael et al. (1989); Car (1993); Brandes (2008) as (see Appendix A),
@@
an expression plotted in Fig. 2(b). The time $\tau$ waited between successive photoelectric emissions detected with perfect efficiency follows the superposition of the quantum beat [term $\propto\cos(\nu\tau)$] onto Rabi oscillations, the latter being developed as a characteristic ringing for increasing drive strength [last term in Eq. (6)]. In contrast to ordinary resonance fluorescence, the distribution minima do not reach zero but are rather bound by the asymptote $\kappa\frac{1}{3}e^{-\kappa\tau}$ defined by the occupation of the intermediate levels in the cascaded process. The higher bound constraining the superposition of oscillations is reached in the limit $\varepsilon_{d}/\kappa\gg 1$. Figure 2(b) shows very good agreement between Eq. (6) and the numerical solution of the ME (1) combined with applying the quantum regression formula.
@@ IV Wave aspect: the spectrum of squeezing
Setting $r=0$ cancels the triggering in the wave-particle correlator and actuates solely the BHD arm, typically used for the measurement of amplitude-squeezing based on the ability of light to interfere with a strong coherent local oscillator (LO) field. These signatures speak of the continuous (wave) character of light. The spectrum of squeezing can be calculated by constructing the photocurrent autocorrelation $\overline{i(t)i(t+\tau)}$ for the selected quadrature of the field $A_{\theta}\equiv\frac{1}{2}(ae^{-i\theta}+a^{\dagger}e^{i\theta})$ as a time average over ergodic records Cresser (2001), and taking the Fourier transform Carmichael (2008).
@@
Let us now revert to unconditional dynamics obeying the master equation (1) with reference to the nonclassical attributes of the particle and wave aspects separately, each revealed by a suitable multi-time correlation function. Figure 4 presents the complementary nonexclusive probability densities expressed through the normalized second-order correlation function Car (1993) for the forward-scattered field. We witness the transition from photon bunching to antibunching as steady-state bimodality builds up alongside a more pronounced Rabi splitting. Consequences of this change in photon statistics are revealed in the individual quantum trajectories of Figs. 5(a ,b), where the two occurring photon emissions are being spaced further apart in the course of an unraveling which combines both particle and wave aspects of light. When we move away from the two-photon resonance peak we come across a superposition of quantum beats with different frequencies and smaller amplitudes [see Fig 4(d)], a consequence of significant detuning of the drive field from an effective two-level resonance structure formed between the dressed JC eigenstates.
@@ V Particles triggering the sampling of waves
At this stage, let us put both arms in operation by setting $r=1/2$ to produce a sampling of a nonlinear stochastic Schrödinger evolution solved by the conditioned wavefunction $|\psi_{c}(t)\rangle$. Our interest is with the full function of the wave-particle correlator, with reference to Fig. 1, in its ability to jointly detect the dual aspects of the incoming radiation. The set $(r,\theta)$ determines each one of the infinite unravelings of the ME: $10<r<$ 1, while the phase $\theta$ of the LO selects the quadrature of interest. The negative spectrum of squeezing we previously depicted for the two-photon resonance driving of the JC oscillator signifies a redistribution of fluctuations between the different quadratures of the cavity field. Here we aim to analyze the quantum fluctuations of the field as ostensible deviations from the steady state, combining conditional measurements with the process of quantum measurement.
We write the master equation for the source – a driven JC oscillator in its zero “system size limit” $\gamma/(2\kappa)\to 0$ – extended now to include the LO mode tuned to the frequency of the coherent drive $\omega_{d}$. The density matrix $\tilde{\rho}$ of the composite system in a frame rotating with $\omega_{d}$ reads Reiner et al. (2001):
@@
where $c$ and $c^{\dagger}$ are the raising and lowering operators for the LO driven mode with excitation amplitude $\beta$. Under the BHD scheme of Fig. 1, the total measured fields at the photodetectors 1 and 2 (in photon flux units) read
@@
Under the assumption that the LO mode is unaffected by the signal mode, the density matrix factorizes into a piece corresponding to the driven JC oscillator under photon blockade, and a piece corresponding to the LO, $\tilde{\rho}=\tilde{\rho}_{s}|\beta\rangle\langle\beta|$. Defining the LO flux as $f\equiv\kappa_{\rm LO}|\beta|^{2}$, we construct the following collapse super-operators,
@@
There are two types of detections which take place under the conditioned balanced homodyne detection (BHD) measurement scheme. The first comes from Eq. (10b) and corresponds to detection “clicks” at the avalanche photodiode (APD). These collapses occur with probability $2\kappa r\langle\psi_{c}(t)|a^{\dagger}a|\psi_{c}(t)\rangle dt$. The second type of collapses concerns the “clicks” registered through the BHD. In any realistic homodyne measurement the LO photon flux $f$ is many orders of magnitude larger than the signal flux, here $2\kappa(1-r)\langle\psi_{c}(t)|a^{\dagger}a|\psi_{c}(t)\rangle$. This means that under the action of (8), a photoelectric emission corresponds with high probability to an annihilation of a LO photon. There is only a small probability ${}f/[2\kappa\langle\psi_{c}(t)|a^{\dagger}a|\psi_{c}(t)\rangle]$ that a photon originated from the JC oscillator. We should note here that these two possibilities exist as a superposition and not as a classical choice, either one or the other. Despite the fact that the collapses are very small, on the characteristic timescale $(2\kappa)^{-1}$ for fluctuations in the signal field they occur very often. The quantum mapping into a stochastic differential equation – a Schrödinger equation with a stochastic non-Hermitian Hamiltonian – involves a coarse graining in time, an expansion of the non-unitary evolution in powers of $\sqrt{2\kappa(1-r)/f}$ and the consideration of a stochastic process for the number of the emitted photoelectrons which depends on a conditioned wavefunction that satisfies the very same Schrödinger equation. Taking that path leads to the following expressions for the difference in photocurrents between detectors 1 and 2 (in Fig. 1), and the wavefunction propagation between photodetections at the APD:
@@
In Eqs. (12) and (13), $\langle A_{\theta}(t)\rangle_{c}$ is the conditioned average of the quadrature amplitude selected by LO, $\langle A_{\theta}(t)\rangle_{c}\equiv\tfrac{1}{2}\langle\psi_{c}(t)|e^{i\theta}a^{\dagger}+ae^{-i\theta}|\psi_{c}(t)\rangle$, calculated with the normalized conditioned wavefunction $|\psi_{c}(t)\rangle$, $B$ is the detection bandwidth and $dW_{t}$ is the Wiener noise increment, the same in both equations (12), (13). In the Monte-Carlo algorithm developed to simulate the wave-particle correlator unravelling, the stochastic differential equation (13) is solved by means of an explicit order 2.0 weak scheme proposed by Kloeden and Platen Kloeden and Platen (1999).
Figure 5: Sampling a homodyne current under low-amplitude nonlinearity. Individual trajectories of the conditioned intracavity photon number (positive-valued plots) and the conditioned field (plots with alternating sign), measured for $r=0.5$ with initial condition $\rho(0)=|1,-\rangle\langle 1,-|$ at different values of the driving frequency and strength $(\Delta\omega_{d}/g,\varepsilon_{d}/g)$: (a) $(-0.7114,0.03)$, (b, c) $(-0.7114,0.055)$ and (d) $(0.545,0.16)$. The inset contour plots in each frame depict the steady-state Wigner function, while the white dashed line indicates the direction set by the phase $\theta$ of the LO. The two additional insets in frame (d) depict the variation of the correlation function of the forward scattered photons $g_{\rm ss}^{(2)}(\tau)$ over an average photon lifetime, at the peaks of the three ($n=3$) and four ($n=4$) photon resonances for $\varepsilon_{d}/g=0.16$. The shaded insets of (b, c) depict unconditional wave-particle correlation functions $h_{\pi/4}(\tau)$ and $h_{3\pi/4}(\tau)$, respectively, normalized to unity at long delays. In all frames, $g/\kappa=200$ and $\gamma/(2\kappa)=0$. A MC algorithm unravelling the ME (1) under the action of the wave-particle correlator was developed ad hoc in Matlab using a basis truncated at the 14-photon level Kloeden and Platen (1999).